What is the H+ concentration for an aqueous soluti

What is the H+ concentration for an aqueous soluti
Answer #1
Concepts and reason

The concept used to solve this problem is based on pH of a solution.

Any substance that increase the concentration of H+{{\rm{H}}^ + } is defined as an acid while a base increase concentration of OH{\rm{O}}{{\rm{H}}^ - } . The concentrations of these ions can vary over a wide range. To avoid dealing these numbers, the concentrations of H+{{\rm{H}}^ + } and OH{\rm{O}}{{\rm{H}}^ - } are reported in terms of pH and pOH.

Fundamentals

The pH of a solution is calculated as follow.

pH=log[H+]{\rm{pH}} = - \log \left[ {{{\rm{H}}^ + }} \right]

Here, [H+]\left[ {{{\rm{H}}^ + }} \right] is the concentration of H+{{\rm{H}}^ + } .

And, pOH is calculated as follow.

pH=log[OH]{\rm{pH}} = - \log \left[ {{\rm{O}}{{\rm{H}}^ - }} \right]

Here, [OH]\left[ {{\rm{O}}{{\rm{H}}^ - }} \right] is concentration of OH{\rm{O}}{{\rm{H}}^ - } .

Part 1

The relation between pH and pOH is as follows:

pH+pOH=14{\rm{pH}} + {\rm{pOH}} = 14

Rearrange this equation and substitute 3.58 for pOH.

pH=14pOH=143.58=10.42\begin{array}{c}\\{\rm{pH}} = 14 - {\rm{pOH}}\\\\{\rm{ = 14}} - {\rm{3}}{\rm{.58}}\\\\{\rm{ = 10}}{\rm{.42}}\\\end{array}

Substitute 10.42 in pH equation.

pH=log[H+]10.42=log[H+][H+]=antilog(10.42)[H+]=38.02×1012M\begin{array}{c}\\{\rm{pH}} = - \log \left[ {{{\rm{H}}^ + }} \right]\\\\10.42 = - \log \left[ {{{\rm{H}}^ + }} \right]\\\\\left[ {{{\rm{H}}^ + }} \right] = {\rm{antilog}}\left( { - 10.42} \right)\\\\\left[ {{{\rm{H}}^ + }} \right] = 38.02 \times {10^{ - 12}}{\rm{ M}}\\\end{array}

Part 2

Calculate the pH from the concentration of HCl in pH equation. Thus, substitute 0.0023 M for H+{{\rm{H}}^ + } .

pH=log[0.0023M]=(2.64)=2.64\begin{array}{c}\\{\rm{pH}} = - \log \left[ {0.0023{\rm{ M}}} \right]\\\\{\rm{ }} = - \left( { - {\rm{2}}{\rm{.64}}} \right)\\\\{\rm{ }} = {\rm{2}}{\rm{.64}}\\\end{array}

Now, the relation between pH and the pOH is as follows:

pH+pOH=14{\rm{pH}} + {\rm{pOH}} = 14

Thus, for third solution rearrange this equation and substitute 8.55 for pOH.

pH=14pOH=148.55=5.45\begin{array}{c}\\{\rm{pH}} = 14 - {\rm{pOH}}\\\\{\rm{ = 14}} - 8.55\\\\{\rm{ = 5}}{\rm{.45}}\\\end{array}

Now, for fourth solution substitute 0.0018 M for OH{\rm{O}}{{\rm{H}}^ - } in pOH equation.

pOH=log[0.0018M]=(2.75)=2.75\begin{array}{c}\\{\rm{pOH}} = - \log \left[ {0.0018{\rm{ M}}} \right]\\\\{\rm{ }} = - \left( { - {\rm{2}}{\rm{.75}}} \right)\\\\{\rm{ }} = {\rm{2}}{\rm{.75}}\\\end{array}

Now, the relation between pH and the pOH is as follows:

pH+pOH=14{\rm{pH}} + {\rm{pOH}} = 14

Rearrange this equation and substitute 2.75 for pOH.

pH=14pOH=142.75=11.25\begin{array}{c}\\{\rm{pH}} = 14 - {\rm{pOH}}\\\\{\rm{ }} = {\rm{14}} - 2.{\rm{75}}\\\\{\rm{ }} = {\rm{11}}{\rm{.25}}\\\end{array}

Now, compare the four values of pH. Thus, the order of four solutions in decreasing acidity is as follows:

1>2=3>41 > 2 = 3 > 4

Part 3

First convert the pH into H+{{\rm{H}}^ + } concentration. Then, substitute 7.21 for pH.

pH=log[H+]7.21=log[H+][H+]=antilog(7.21)[H+]=61.80×109M\begin{array}{c}\\{\rm{pH}} = - \log \left[ {{{\rm{H}}^ + }} \right]\\\\7.21 = - \log \left[ {{{\rm{H}}^ + }} \right]\\\\\left[ {{{\rm{H}}^ + }} \right] = {\rm{antilog}}\left( { - 7.21} \right)\\\\\left[ {{{\rm{H}}^ + }} \right] = 61.80 \times {10^{ - 9}}{\rm{ M}}\\\end{array}

Since, the solution is neutral thus,

pH=pOH{\rm{pH}} = {\rm{pOH}}

Therefore, the ionic product of water is written as follows:

Kw=[H+][OH]=[H+]2\begin{array}{c}\\{K_w} = \left[ {{{\rm{H}}^ + }} \right]\left[ {{\rm{O}}{{\rm{H}}^ - }} \right]\\\\ = {\left[ {{{\rm{H}}^ + }} \right]^2}\\\end{array}

Substitute 61.80×109M61.80 \times {10^{ - 9}}{\rm{ M}} for [H+]\left[ {{{\rm{H}}^ + }} \right] .

Kw=[61.8×109]2=3.8×1015\begin{array}{c}\\{K_w} = {\left[ {61.8 \times {{10}^{ - 9}}} \right]^2}\\\\{\rm{ = }}3.8 \times {10^{ - 15}}\\\end{array}

Ans: Part 1

The concentration of H+{{\rm{H}}^ + } is 38.02×1012M38.02 \times {10^{ - 12}}{\rm{ M}} .

Part 2

The order of four solutions in decreasing acidity is as follows:

1>2=3>41 > 2 = 3 > 4

Part 3

The ionic product of water is equal to 3.8×10153.8 \times {10^{ - 15}} .